OpenStudy (anonymous):

Given a = -3, b = 4 and c = -5, evaluate a - b - c.

OpenStudy (the_fizicx99):

a = -3 b = 4 c = -5 -3 - 4 - (-5) Solve?

OpenStudy (anonymous):

how

OpenStudy (the_fizicx99):

-3 -4 = ?

OpenStudy (anonymous):

what is it

OpenStudy (the_fizicx99):

._. I'm not giving you the answer. Solve it, -3 - 4= ??

OpenStudy (anonymous):

1

OpenStudy (the_fizicx99):

No, 3 is negative, -3 - 4 ?